# On sporadic symmetry breaking operators for principal series representations of the de Sitter and Lorentz groups

### v1

@@ Abstract
In this paper, we construct and classify all differential symmetry breaking operators $\mathbb{D}_{\lambda,\nu}^{N,m}:C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1})\rightarrow C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$ between principal series representations with respect to the restriction $SO_{0}(4,1)\downarrow SO_{0}(3,1)$ when the parameters satisfy the condition $|m|>N$. In this case we also prove a localness theorem, namely, all symmetry breaking operators between the principal series representations above are necessarily differential operators. In addition, we show that all these symmetry breaking operators are sporadic in the sense of T. Kobayashi, that is, they cannot be obtained by residue formulas of distributional kernels.
@@ 1 Introduction
The problem of constructing and classifying all SBOs given a concrete pair $(G,G^{\prime})$ is a challenging one. It was first fully solved in 2018 for the pair $(G,G^{\prime})=(O(n+1,1),O(n,1))$ in the case of differential forms. Concretely, all SBOs that can be written as differential operators—referred to as differential symmetry breaking operators (DSBOs)—were constructed and classified in [20], while the non-local ones were addressed in [27, 28].
The family of DSBOs constitutes a proper and significant subfamily within the set of all SBOs. In particular, there exists a duality between the space of DSBOs and the space of certain homomorphisms between generalized Verma modules (see [25, Thm. 2.9]), highlighting the fact that the construction and classification of DSBOs is also algebraically meaningful. Moreover, in some special cases, the sets of SBOs and DSBOs coincide—a phenomenon known as localness theorem (cf. [25, Thm. 5.3], [28, Thm. 3.6])—thus reducing the problem of constructing SBOs to that of constructing only the differential ones. For these reasons, among others, substantial progress has been made in recent years toward the construction and classification of DSBOs.
@@
Following this line of study of symmetry breaking operators, we turn our attention to the connected pair $(G,G^{\prime})=(SO_{0}(n+1,1),SO_{0}(n,1))$ and aim to construct and classify all SBOs. For certain principal series representations, this problem is expected to share strong connections with the corresponding problem for the non-connected pair $(O(n+1,1),O(n,1))$, and it is plausible that some of the techniques developed in [27, 28] may be adapted to the connected case. However, the problem for general principal series representations remains open, even for $n=3$.
@@
Here, the spaces $C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1})$ and $C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$ stand for general principal series representations of $SO_{0}(4,1)$ and $SO_{0}(3,1)$ respectively (for a concrete definition, see Section 2).
@@
is non-zero. In particular, determine
@@
We give a complete solution to these two problems for $|m|>N$ in Theorems 1.3 and 1.4, respectively.
Although this paper primarily focuses on the case $|m|>N$, Problems A and B have been addressed for some other cases in the past. For example, when $m=N=0$, a solution was obtained from a conformal geometric perspective in [15, 23]. For $N=1$ and $m=0$, a solution can be deduced from the results of [20], while the case $|m|\geq N=1$ was completely solved in [35]. Furthermore, for general $N\in{\mathbb{N}}$, the case $|m|=N$ is treated in [36]. A summary of the known and open cases is presented in Table 1.1 below.
Although Problems A and B above remain unsolved for $0\leq|m|<N$, the case $|m|>N$ treated in this work corresponds to the hardest one, as the SBOs appearing are all sporadic in the sense of Kobayashi–Speh (see Definition 7.1). That is, they cannot be obtained as residues of distributional kernels, contrarily to the regular SBOs (cf. [27, 28]). We prove this fact in Section 7.
@@ 1.1 Organization of the paper
This paper is structured into eight sections, including the Introduction and an Appendix. In Section 1.2 we present our main results, which provide a complete solution to Problems 1.1 and 1.2 in the case $|m|>N$. In particular, Theorems 1.3 and 1.4 address Problem 1.2, and Theorem 1.6 solves Problem 1.1 by establishing a localness theorem that reduces it to Problem 1.2.
Section 2 provides a concrete geometric description of the setting. In Section 3, we review and apply the F-method, which serves as the primary tool for solving Problem 1.2. This method allows us to reduce Problem 1.2 to solving a specific system of ordinary differential equations, as formulated in Theorems 3.5 and 3.6.
We show in Section 5 that it suffices to consider the case $m>N$, since the case $m<-N$ follows from the former one (see Proposition 5.1). The proofs of Theorems 1.3 and 1.4 are given in Section 3.2 by using Theorems 3.5 and 3.6, while the proof of Theorem 3.6—the most technically demanding part of this work—is presented in Section 4. (We note that a proof of Theorem 3.5 has already been given in [36].)
@@ 1.2 Main results
In this subsection we state our main results, providing a complete solution to Problems A and B for $|m|>N$. We also introduce a localness theorem, which allows us to solve not only Problem 1.2, but also Problem 1.1 in the same parameter range. We start with the result that solves Problem A for $|m|>N$.
@@
Suppose $\nu-\lambda\in{\mathbb{N}}$, and let ~$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ denote the following (scalar-valued) differential operator, sometimes called the Juhl operator (cf. [20, 2.22]):
~$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ $=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-4\frac{\partial^{2}}{\partial z\partial\overline{z}},\frac{\partial}{\partial x_{3}}\right),$ (1.3)
@@
$\Gamma(d,r)$ $:=\begin{pmatrix}d\\ r\end{pmatrix}\frac{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|+N-d-1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|-1}{2}\right]\right)}\frac{\Gamma(2N-r+1)}{\Gamma(N+1)}\frac{\Gamma(N+|m|+1)}{\Gamma(N+|m|+1-r)},$ (1.4)
$A(d,r)$ $:=(-1)^{\nu-1}\Gamma(d,r)\frac{\Gamma\left(\lambda+\left[\frac{-\nu-\lambda-|m|+N-d+1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|+N-d-1}{2}\right]\right)}\frac{\Gamma(N+\nu)}{\Gamma(N+\nu-r)},$
$B(d,r)$ $:=(-1)^{d-N}\Gamma(2N-d,r)\frac{\Gamma(N-\nu+2)}{\Gamma(N-\nu+2-r)}.$
@@
Let $N\in{\mathbb{N}}$ and $m\in{\mathbb{Z}}$ with $|m|>N$, and let $\lambda,\nu\in{\mathbb{C}}$ with $\lambda\in{\mathbb{Z}}_{\leq 1-|m|}$ and $\nu\in[1-N,N+1]\cap{\mathbb{Z}}$. Then, any differential symmetry breaking operator in (1.2) is proportional to the differential operator $\mathbb{D}_{\lambda,\nu}^{N,m}$ of order $\nu-\lambda$ given as follows:
@@
Both Theorems 1.3 and 1.4 will be proved in Section 3.2.
Remark 1.5.
The case $|m|>N=1$ was solved in [35], where the generator of (1.2) is given as follows for $\lambda\in{\mathbb{Z}}_{\leq 1-|m|}$ and $\nu=0,1,2$.
[…]
$2^{2\nu}(-1)^{\nu+1}\frac{\Gamma\left(\lambda+\left[\frac{-\lambda-\nu-m+2}{2}\right]\right)}{\Gamma(\lambda+\left[\frac{\nu-\lambda-m-1}{2}\right])}\widetilde{{\mathbb{C}}}_{\lambda+1,2-\nu-m}\frac{\partial^{2\nu+m-1}}{\partial z^{\nu}\partial\overline{z}^{\nu+m-1}}\otimes u_{0}^{\vee}\\ +\left((m(1-\nu)-\lambda+2)\widetilde{{\mathbb{C}}}_{\lambda,\nu-m}-2\gamma(\lambda-1,\nu-\lambda-m)\frac{\partial}{\partial x_{3}}\widetilde{{\mathbb{C}}}_{\lambda+1,\nu-m}\right)\frac{\partial^{m}}{\partial\overline{z}^{m}}\otimes u_{1}^{\vee}\\ -4\gamma(\lambda-1,\nu-\lambda-m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu-m}\frac{\partial^{m+1}}{\partial\overline{z}^{m+1}}\otimes u_{2}^{\vee},$ (1.8)
$-4\gamma(\lambda-1,\nu-\lambda+m)\widetilde{{\mathbb{C}}}_{\lambda+1,\nu+m}\frac{\partial^{-m+1}}{\partial z^{-m+1}}\otimes u_{0}^{\vee}\\ +\left((m(1-\nu)+\lambda-2)\widetilde{{\mathbb{C}}}_{\lambda,\nu+m}+2\gamma(\lambda-1,\nu-\lambda+m)\frac{\partial}{\partial x_{3}}\widetilde{{\mathbb{C}}}_{\lambda+1,\nu+m}\right)\frac{\partial^{-m}}{\partial z^{-m}}\otimes u_{1}^{\vee}\\ +2^{2\nu}(-1)^{\nu+1}\frac{\Gamma\left(\lambda+\left[\frac{-\lambda-\nu+m+2}{2}\right]\right)}{\Gamma(\lambda+\left[\frac{\nu-\lambda+m-1}{2}\right])}\widetilde{{\mathbb{C}}}_{\lambda+1,2-\nu+m}\frac{\partial^{2\nu-m-1}}{\partial z^{\nu-m-1}\partial\overline{z}^{\nu}}\otimes u_{2}^{\vee}.$ (1.9)
Here, $\gamma(\mu,\ell)$ denotes the constant defined in (8.11).
$\mathbb{D}_{\lambda,\nu}^{m}=2\mathbb{D}_{\lambda,\nu}^{1,m}.$ (1.10)
In fact, by (1.5), $\mathbb{D}_{\lambda,\nu}^{1,m}$ amounts to
On the other hand, from the three-term relation [35, Eq. (9.5)], we have
[…]
And by (1.4), the following holds:
$2B(1,0)$ $=4,$
$\frac{1}{2}B(1,1)$ $=\frac{1}{2}(m+1)(2-\nu),$
$B(2,0)$ $=-2\gamma(\lambda-1,\nu-\lambda-m).$
Now, (1.10) clearly follows.
@@ 2 Review of Principal Series Representations of the de Sitter and Lorentz Groups
$O(4,1)=\{g\in GL(5,{\mathbb{R}}):Q_{4,1}(gx)=Q_{4,1}(x)\enspace\forall x\in{\mathbb{R}}^{5}\},$
@@
$H_{0}$ $:=E_{1,5}+E_{5,1}$ (2.1)
$N_{j}^{+}$ $:=-E_{1,j+1}+E_{j+1,1}-E_{5,j+1}-E_{j+1,5}$ $(1\leq j\leq 3),$
$N_{j}^{-}$ $:=-E_{1,j+1}+E_{j+1,1}+E_{5,j+1}+E_{j+1,5}$ $(1\leq j\leq 3),$
$X_{p,q}$ $:=E_{q+1,p+1}-E_{p+1,q+1}$ $(1\leq p\leq q\leq 3).$
Observe that the sets $\{N_{j}^{+}:1\leq j\leq 3\}$, $\{N_{j}^{-}:1\leq j\leq 3\}$ and $\{X_{p,q}:1\leq p,q\leq 3\}\cup\{H_{0}\}$ determine bases of the following Lie subalgebras of $\mathfrak{g}({\mathbb{R}})$:
$\mathfrak{n}_{+}({\mathbb{R}})$ $:=\ker(\text{{ad}}(H_{0})-\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{n}_{-}({\mathbb{R}})$ $:=\ker(\text{{ad}}(H_{0})+\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{l}({\mathbb{R}})$ $:=\ker\text{{ad}}(H_{0})\subset\mathfrak{g}({\mathbb{R}}).$
@@
where the second map comes from the projection $\Xi\rightarrow S^{3},\enspace x\mapsto\frac{1}{x_{4}}{}^{t}(x_{0},x_{1},x_{2},x_{3})$. The group $P\subset G$ is a parabolic subgroup of $G$ with Langlands decomposition as follows:
@@
$\begin{gathered}n_{-}(x):=\exp\left(\sum_{j=1}^{3}x_{j}N_{j}^{-}\right)=I_{5}+\begin{pmatrix}-\frac{1}{2}\|x\|^{2}&-{}^{t}x&-\frac{1}{2}\|x\|^{2})\\ x&0&x\\ \frac{1}{2}\|x\|^{2}&{}^{t}x&\frac{1}{2}\|x\|^{2}\end{pmatrix},\end{gathered}$ (2.2)
@@
Moreover, for any $N\in{\mathbb{N}}$, let $(\sigma^{2N+1},V^{2N+1})$ be the $(2N+1)$-dimensional irreducible representation of $M\simeq SO(3)$ realized in $V^{2N+1}:=\text{{Pol}}^{2N}({\mathbb{C}}^{2})$, the space of homogeneous polynomials of degree $2N$ in two variables (for details about this realization, see [36]). Now, we consider the representation of $P=MAN_{+}$ defined as $\sigma_{\lambda}^{2N+1}:=\sigma^{2N+1}\boxtimes{\mathbb{C}}_{\lambda}\boxtimes\text{{id}}$ and define the homogeneous vector bundle over $G/P\simeq S^{3}$ associated to $(\sigma_{\lambda}^{2N+1},V_{\lambda}^{2N+1}):$
@@
Now, for any $\nu\in{\mathbb{C}}$ and any $m\in{\mathbb{Z}}$, we consider the representation of $P^{\prime}$ defined as $\tau_{m,\nu}:={\mathbb{C}}_{m}\boxtimes{\mathbb{C}}_{\nu}\boxtimes\text{{id}}$, and define the homogeneous vector bundle over $G^{\prime}/P^{\prime}\simeq S^{2}$ associated to $(\tau_{m,\nu},{\mathbb{C}}_{m,\nu}):$

### v2

@@ Abstract
In this paper, we construct and classify all differential symmetry breaking operators between certain principal series representations of the pair $SO_{0}(4,1)\supset SO_{0}(3,1)$. In this case, we also prove a localness theorem, namely, all symmetry breaking operators between the principal series representations in concern are necessarily differential operators. In addition, we show that all these symmetry breaking operators are sporadic in the sense of T. Kobayashi, that is, they cannot be obtained by residue formulas of meromorphic families of symmetry breaking operators.
@@ 1 Introduction
The problem of constructing and classifying all SBOs given a concrete pair $(G,G^{\prime})$ is a challenging one. One of the first results was obtained for the pair $(G,G^{\prime})=(O(n+1,1),O(n,1))$ in the case of principal series representations induced from exterior power representations. Concretely, all SBOs that can be written as differential operators—referred to as differential symmetry breaking operators (DSBOs)—were constructed and classified in [23], while the non-local ones were addressed in [30, 31]. Another early result along these lines was given for the pair $(G,G^{\prime})=(SO_{0}(1,n)\times SO_{0}(1,n),\text{{diag}}(SO_{0}(1,n)))$ for spherical principal series representations ([5, 6]).
The family of DSBOs constitutes a proper and significant subfamily within the set of all SBOs. In particular, there exists a duality between the space of DSBOs and the space of certain homomorphisms between generalized Verma modules (see [28, Thm. 2.9]), highlighting the fact that the construction and classification of DSBOs is also algebraically meaningful. Moreover, in some special cases, the sets of SBOs and DSBOs coincide—a phenomenon known as localness theorem (cf. [28, Thm. 5.3], [31, Thm. 3.6])—thus reducing the problem of constructing SBOs to that of constructing only the differential ones. For these reasons, among others, substantial progress has been made in recent years toward the construction and classification of DSBOs.
@@
Following this line of study of symmetry breaking operators, we turn our attention to the connected pair $(G,G^{\prime})=(SO_{0}(n+1,1),SO_{0}(n,1))$ and aim to construct and classify all SBOs. For certain principal series representations, this problem shares strong connections with the corresponding problem for the non-connected pair $(O(n+1,1),O(n,1))$, and some of the techniques developed in [30, 31] can be adapted to the connected case. However, the problem for general principal series representations remains open, even for $n=3$ (see, for instance Table 1.1).
@@
Here, $C^{\infty}(S^{3},\mathcal{V}_{\lambda}^{2N+1})$ and $C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$ denote principal series representations induced from minimal parabolic subgroups of $SO_{0}(4,1)$ and $SO_{0}(3,1)$, respectively. They are realized as spaces of smooth sections over the flag varieties $S^{3}$ and $S^{2}$. The former is parametrized by a natural number $N\in{\mathbb{N}}$, corresponding to the induced representation of the compact factor of the Levi subgroup of dimension $2N+1$, and by a complex parameter $\lambda\in{\mathbb{C}}$. The latter is parametrized by an integer number $m\in{\mathbb{Z}}$, corresponding to the characters of the compact factor of the Levi subgroup, and by a complex number $\nu\in{\mathbb{C}}$ (these are precisely defined in Section 2).
@@
is non-zero. Moreover, determine
@@
We give a complete solution to these two problems for $|m|>N$ in Theorems 1.3 and 1.5, respectively.
Although this paper primarily focuses on the case $|m|>N$, Problems A and B have been addressed for some other cases in the past. For example, when $m=N=0$, a solution was obtained from a conformal geometric perspective in [18, 26]. For $N=1$ and $m=0$, a solution can be deduced from the results of [23], while the case $|m|\geq N=1$ was completely solved in [38]. Furthermore, for general $N\in{\mathbb{N}}$, the case $|m|=N$ is treated in [39]. A summary of the known and open cases is presented in Table 1.1.
Although Problems A and B above remain unsolved for $0\leq|m|<N$, the case $|m|>N$ treated in this work corresponds to the hardest one, as the SBOs appearing are all sporadic in the sense of T. Kobayashi (see Definition 7.1). That is, they cannot be obtained as residues of a meromorphic family of SBOs, contrarily to the regular SBOs (cf. [30, 31]). We prove this fact in Section 7.
@@ 1.1 Organization of the paper
This paper is structured into eight sections, including the Introduction and an Appendix. In Section 1.2 we present our main results, which provide a complete solution to Problems 1.1 and 1.2 in the case $|m|>N$. In particular, Theorems 1.3 and 1.5 address Problem 1.2, and Theorem 1.6 solves Problem 1.1 by establishing a localness theorem that reduces it to Problem 1.2.
Section 2 provides a concrete geometric description of the setting. In Section 3, we review and apply the F-method, which serves as the primary tool for solving Problem 1.2. This method allows us to reduce Problem 1.2 to solving a specific system of ordinary differential equations, as formulated in Theorems 3.6 and 3.7.
We show in Section 5 that it suffices to consider the case $m>N$, since the case m<Nm<-N follows from the former one (see Proposition 5.1). The proofs of Theorems 1.3 and 1.5 are given in Section 3.2 by using Theorems 3.6 and 3.7, while the proof of Theorem 3.7—the most technically demanding part of this work—is presented in Section 4. (We note that a proof of Theorem 3.6 has already been given in [39].)
@@ 1.2 Main results
In this subsection we state our main results. We start with the result that solves Problem A for $|m|>N$.
@@
As a consequence of Theorem 1.3 above, we have the following corollary.
Corollary 1.4.
Let $\lambda,\nu\in{\mathbb{C}}$, $N\in{\mathbb{N}}$ and $m\in{\mathbb{Z}}$ with $|m|>N$, and let $\mathbb{D}$ be any non-zero differential symmetry breaking operator in (1.2). Then, the following holds.
  1. (1)
    The principal series representation $C^{\infty}(S^{3},\mathcal{V}^{2N+1}_{\lambda})$ of $SO_{0}(4,1)$ is reducible.
  2. (2)
    The principal series representation $C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$ of $SO_{0}(3,1)$ is irreducible. In particular, $\text{{im}}(\mathbb{D})$ is dense in $C^{\infty}(S^{2},\mathcal{L}_{m,\nu})$.
@@
Suppose $\nu-\lambda\in{\mathbb{N}}$, and let $\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ denote the following (scalar-valued) differential operator, sometimes called the Juhl operator (cf. [23, 2.22]):
$\widetilde{{\mathbb{C}}}_{\lambda,\nu}$ $=\text{{Rest}}_{x_{3}=0}\circ\left(I_{\nu-\lambda}\widetilde{C}^{\lambda-1}_{\nu-\lambda}\right)\left(-4\frac{\partial^{2}}{\partial z\partial\overline{z}},\frac{\partial}{\partial x_{3}}\right),$ (1.3)
@@
$\Gamma(d,r)$ $=\begin{pmatrix}d\\ r\end{pmatrix}\frac{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|+N-d-1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|-1}{2}\right]\right)}\frac{\Gamma(2N-r+1)}{\Gamma(N+1)}\frac{\Gamma(N+|m|+1)}{\Gamma(N+|m|+1-r)},$ (1.4)
$A(d,r)$ $=(-1)^{\nu-1}\Gamma(d,r)\frac{\Gamma\left(\lambda+\left[\frac{-\nu-\lambda-|m|+N-d+1}{2}\right]\right)}{\Gamma\left(\lambda+\left[\frac{\nu-\lambda-|m|+N-d-1}{2}\right]\right)}\frac{\Gamma(N+\nu)}{\Gamma(N+\nu-r)},$
$B(d,r)$ $=(-1)^{d-N}\Gamma(2N-d,r)\frac{\Gamma(N-\nu+2)}{\Gamma(N-\nu+2-r)}.$
@@
Let $N\in{\mathbb{N}}$ and $m\in{\mathbb{Z}}$ with $|m|>N$, and let $\lambda\in{\mathbb{Z}}_{\leq 1-|m|}$ and $\nu\in[1-N,N+1]\cap{\mathbb{Z}}$. Then, any differential symmetry breaking operator in (1.2) is proportional to the differential operator $\mathbb{D}_{\lambda,\nu}^{N,m}$ of order $\nu-\lambda$ given as follows:
@@
Both Theorems 1.3 and 1.5 will be proved in Section 3.2, while Corollary 1.4 will be proved in Section 3.3
@@ 2 Review of Principal Series Representations of the de Sitter and Lorentz Groups
$O(4,1)=\{g\in GL(5,{\mathbb{R}}):Q_{4,1}(gx)=Q_{4,1}(x)\kern 4.62497pt\forall x\in{\mathbb{R}}^{5}\},$
@@
$H_{0}$ $=E_{1,5}+E_{5,1}$ (2.1)
$N_{j}^{+}$ $=-E_{1,j+1}+E_{j+1,1}-E_{5,j+1}-E_{j+1,5}$ $(1\leq j\leq 3),$
$N_{j}^{-}$ $=-E_{1,j+1}+E_{j+1,1}+E_{5,j+1}+E_{j+1,5}$ $(1\leq j\leq 3),$
$X_{p,q}$ $=E_{q+1,p+1}-E_{p+1,q+1}$ $(1\leq p\leq q\leq 3).$
Observe that the sets $\{N_{j}^{+}:1\leq j\leq 3\}$, $\{N_{j}^{-}:1\leq j\leq 3\}$ and $\{X_{p,q}:1\leq p,q\leq 3\}\cup\{H_{0}\}$ form bases of the following Lie subalgebras of $\mathfrak{g}({\mathbb{R}})$:
$\mathfrak{n}_{+}({\mathbb{R}})$ $=\ker(\text{{ad}}(H_{0})-\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{n}_{-}({\mathbb{R}})$ $=\ker(\text{{ad}}(H_{0})+\text{{id}})\subset\mathfrak{g}({\mathbb{R}}),$
$\mathfrak{l}({\mathbb{R}})$ $=\ker\text{{ad}}(H_{0})\subset\mathfrak{g}({\mathbb{R}}).$
@@
[…]
@@
[…]
@@
Moreover, for any $N\in{\mathbb{N}}$, let $(\sigma^{2N+1},V^{2N+1})$ be the $(2N+1)$-dimensional irreducible representation of $M\simeq SO(3)$ realized in $V^{2N+1}:=\text{{Pol}}^{2N}({\mathbb{C}}^{2})$, the space of homogeneous polynomials of degree $2N$ in two variables (for details about this realization, see [39]). Now, we consider the representation of $P=MAN_{+}$ defined as $\sigma_{\lambda}^{2N+1}:=\sigma^{2N+1}\boxtimes{\mathbb{C}}_{\lambda}\boxtimes\mathbf{1}$ and define the homogeneous vector bundle over $G/P\simeq S^{3}$ associated to $(\sigma_{\lambda}^{2N+1},V_{\lambda}^{2N+1}):$
@@
Now, for any $\nu\in{\mathbb{C}}$ and any $m\in{\mathbb{Z}}$, we consider the representation of $P^{\prime}$ defined as $\tau_{m,\nu}:={\mathbb{C}}_{m}\boxtimes{\mathbb{C}}_{\nu}\boxtimes\mathbf{1}$, and define the homogeneous vector bundle over $G^{\prime}/P^{\prime}\simeq S^{2}$ associated to $(\tau_{m,\nu},{\mathbb{C}}_{m,\nu}):$